How many pounds of hamburger that costs $1.10 per pound must be mixed with 50 pounds of hamburger that costs $1.80 per pound to make a mixture that costs $1.30 per pound?
Answer by ankor@dixie-net.com(12701)

How many pounds of hamburger that costs $1.10 per pound must be mixed with 50 pounds of hamburger that costs $1.80 per pound to make a mixture that costs $1.30 per pound?
:
Let x = number of lbs of $1.10 hamburger
:
We know that there is 50 lbs of the $1.80 hamburger
:
Therefore the resulting amt will be (x + 50)
:
The equation:
1.1x + 1.8(50) = 1.3(x + 50)
:
1.1x + 90 = 1.3x + 65
:
1.1x - 1.3x = 65 - 90
:
-.2x = -25
:
x = -25/-.2
:
x = +125 lbs of $1.10 stuff required
:
:
Check:
1.1(125) + 1.8(50) = 1.3(125 + 50)
137.5 + 90 = 1.3(175)
227.5 = 227.5 proves our solution